Adv. Math. Econ. 8, 257-271 (2006)

Implementation with unknown endowments in a two-trader pure exchange economy

Leonid Hurwicz

University of Minesota Received: November 3, 2005 Revised: November 3, 2005 JEL classification: D02, D51, D60, D61 Mathematical Subject Classification (2000): 91B32, 91B50, 91B60